# RationalVLA: A Rational Vision-Language-Action Model with Dual System

### v1

@@ Abstract
A fundamental requirement for real-world robotic deployment is the ability to understand and respond to natural language instructions. Existing language-conditioned manipulation tasks typically assume that instructions are perfectly aligned with the environment. This assumption limits robustness and generalization in realistic scenarios where instructions may be ambiguous, irrelevant, or infeasible. To address this problem, we introduce RAtional MAnipulation (RAMA), a new benchmark that challenges models with both unseen executable instructions and defective ones that should be rejected. In RAMA, we construct a dataset with over 14,000 samples, including diverse defective instructions spanning six dimensions: visual, physical, semantic, motion, safety, and out-of-context. We further propose the Rational Vision-Language-Action model (RationalVLA). It is a dual system for robotic arms that integrates the high-level vision-language model with the low-level manipulation policy by introducing learnable latent space embeddings. This design enables RationalVLA to reason over instructions, reject infeasible commands, and execute manipulation effectively. Experiments demonstrate that RationalVLA outperforms state-of-the-art baselines on RAMA by a 14.5% higher success rate and 0.94 average task length, while maintaining competitive performance on standard manipulation tasks. Real-world trials further validate its effectiveness and robustness in practical applications. Our project page is https://irpn-eai.github.io/rationalvla.
Index Terms:
Robot Manipulation, Vision-Language-Action Model, Dual System
@@ I Introduction
Embodied intelligence is considered as the ultimate form of artificial intelligence (AI) [1]. A necessary condition for the successful deployment of embodied intelligence in the real world is their ability to understand natural language and respond appropriately, either by providing correct answers or by executing the corresponding actions. This demand has sparked research on language-conditioned manipulation tasks, which require robots to follow natural language instructions to complete specific manipulation actions. Recent works have identified and investigated several key challenges in this field, particularly in handling complex and previously unseen instructions [2, 3, 4]. They primarily leverage the language generalization capabilities of pretrained vision-language models (VLMs) [5, 6, 7, 8, 9, 10], fine-tuning them on robotic datasets to obtain vision-language-action (VLA) models [11, 2, 12, 13, 14, 15], which enables the handling of various instructions to some extent.
@@ II Related Work
While previous research[21, 22, 23, 24, 25] has made notable progress in enabling robots to perform precise and efficient manipulation, our work specifically focuses on the challenges of language-conditioned manipulation. To enhance the robustness of instruction understanding in such robots, prior research relevant to our proposed RationalVLA model can be organized into three key areas: (1) language-conditioned manipulation, which focuses on learning action policies from natural language. (2) dual-system VLA models, which decouple high-level reasoning from low-level control. and (3) efforts addressing defective instructions in human-robot collaboration of RAMA benchmark. Below, we briefly review representative work in each area.
@@ II-A Language-conditioned Manipulation
Early works [26, 27, 11] use text encoders to map language instructions to latent features, which are then conditioned by robot policies to predict actions. Some approaches [28, 29, 16, 30] employ diffusion models as action policies. RT-2 [2] and subsequent works [31, 32] integrate large language models to enhance language generalization. Inner Monologue [33] enriches instruction processing and planning in robotic control. RoboFlamingo [3] fine-tunes vision-language models for action prediction, while GR-1 [34] learns from action videos and Vision Language Planning [35] generate futural action videos from language to predict action. RoboMamba [36] utilizes state space models for efficient instruction-based reasoning. VLAS [37] extends the text language to speech. Our model demonstrates superior reasoning capabilities over these methods, enabled by its effective understanding of the intricate relationship between language and the environment. Other work [38] is a dual-arm robot framework combining dynamic manipulation primitives efficiently unfolds garments using self-supervised learning, achieving high coverage without expert data.
@@ II-C Defective Instructions in Human-robot Interaction
Ob-VLN [47] first considers the defective instructions in vision-language navigation. The discrepancies in actual scenes and given instructions can cause major failures. This work constructs a virtual graph to help the robot manage it. BadNaver [48] further identifies defective instructions as a type of jailbreak attack, which leads the model to execute vulnerable actions. Experiments on various robot agents imply that these instructions affect the safety. These works emphasize the importance of dealing with defective instructions. However, these works limit the question in navigation tasks, overlooking its importance in manipulation tasks.
@@ III-A Task Setting
  • Physical. The physical non-executable task refers to the target with non-existent shape and size in the scenario. The understanding of these spatial positions sets requirements for physical prior knowledge of the robot assistant.
@@
  • Motion. The motion non-executable task refers to actions that can not be executed due to the limitations of the structure and kinematic characteristics of the mechanical arm. This requires the robot assistant to assess the feasibility of executing the action based on its capabilities.
@@
  • Out-of-context. In addition to the above dimensions, some completely unrelated, illogical, and nonsensical commands should also be included as non-executable instructions to ensure the robustness robot assistant against noise.
@@ III-B Simulation Environment
To establish a unified and convenient benchmark, we construct a dataset and evaluation protocol based on CALVIN, which is well-regarded in language-conditioned manipulation research. 1) It ensures fair comparisons between different models and reproducibility of experiments. 2) It supports long-horizon tasks, which is crucial for evaluating the influence of defective instructions.
CALVIN is built on top of the PyBullet [51] simulator. It involves a desk with some objects and a Franka Panda Robot arm that manipulates the scene. It consists of 34 tasks and 4 different environments (A, B, C and D).
Figure 2: In RAMA benchmark, we assess models across six dimensions: visual, physical, semantic, motion, safety, and out-of-context.
@@ III-C Dataset
We then convert the language annotations into the chat-like expression of MLLM assistants. CALVIN contains one language instruction and a list of (observation, action) pairs $[x_{task},(o_{0},a_{0},o_{1},a_{1},...,o_{t},a_{t},...)]$ per trajectory. We programmatically generate texts in the format of chat interactions using templates. A simple example of user-assistant interactions, is “User: Can you help me $x_{task}$? Assistant: Sure, I will $x_{task}$ <ACT>." For defective instructions, it will be “User: Can you help me $x_{task}$? Assistant: Sorry, I can not $x_{task}$ <REJ>." This process trains the model to recognize and appropriately respond to everyday human instructions, making informed decisions on whether to execute or refrain from executing actions. It thereby fosters a conversational interface capable of seamlessly transitioning from dialogue to actionable responses.
@@ III-D Evaluation
Figure 5: Overview of RationalVLA. Given the observation and text instruction including the defective one, an MLLM generates a text description of a task and an <ACT> token or a <REJ> token. The latent space embeddings from the <ACT> token’s last layer serve as a high-level latent goal for the downstream policy and those of <REJ> token reject defective instructions in the current scene. Our dual-system model can run the high-level MLLM reasoning and low-level action policy execution loops asynchronously. During inference, the action policy frequently updates actions based on environment changes and the latest <ACT> token’s embedding, while the MLLM updates less frequently, enabling efficient inference.
@@ IV-A Comparison of Architectures
Fine-tuned VLA. Given the strong generalization capabilities on visual and language modalities demonstrated by pretrained VLAs, a natural approach is to fine-tune an end-to-end VLA model to process and understand complex instructions. However, several key challenges limit its performance: (1) Fine-tuning a VLA model often leads to catastrophic forgetting [18], resulting in degraded language capabilities, e.g., RoboFlamingo reaches a mediocre performance on the enriched instructions. (2) Considering downstream robotic tasks, VLA models frequently underperform compared to specialized expert policies, e.g., RoboFlamingo lags behind 3DDA on standard CALVIN benchmarks.
VLM as Identifier. To address the aforementioned issues, a straightforward idea is to leverage both the multimodal reasoning capabilities of pretrained VLMs and the manipulation proficiency of a task-specific policy. Specifically, a separately trained VLM can serve as a high-level identifier that decides whether to reject a given instruction, after which a downstream policy is responsible for task execution. This constitutes a rudimentary dual-system architecture. With the capabilities of powerful VLMs such as GPT-4o, this design can effectively filter out inappropriate instructions. However, such a naïve combination fails to fully exploit the VLM’s language understanding capabilities. When faced with complex instructions, the downstream policy often lacks sufficient comprehension, leading to task failure. Consequently, this approach falls short of achieving our overarching goal: aligning the robotic agent’s action space with the complexity of natural language instructions.
@@ IV-B Our Dual-system VLA
We expand the vocabulary of the language model with an additional <ACT> token and a <REJ> token. The <ACT> token extracts the vision-language information, representing the latent space embedding of the action entity. Upon generating <ACT> token, its hidden embedding is further fed into the low-level policy as a condition. Some previous works have focused on dynamically processing neural network inputs [19, 20], which is equally effective for robots to flexibly handle language instructions. The model is trained to output either <ACT> tokens $y_{<ACT>}$ or <REJ> tokens $y_{<REJ>}$ depending on whether the instructions are executable or not in the current scene. Then, the $y_{<ACT>}$ is projected by a projector $\varPsi$ and further unsqueezed in the zero dimension to align with the policy latent conditioning space. Thus, we get the latent embeddings $z=unsqueenze(\varPsi(y_{<ACT>}))$. For $y_{<REJ>}$, it is projected to be $z_{0}=\varPsi(y_{<REJ>})$, which is an all-zero tensor to stop action. Finally, $z$ is fed into the policy $\pi_{\theta}$.
@@ IV-C Training
The training of RationalVLA employs a combination of techniques to integrate the MLLM and policy components. We leverage Low Rank Adaptation [52] (LoRA) for the supervised fine-tuning of the MLLM, allowing for more efficient training.
Stage 1: Pre-training for feature alignment. We adopt a cold start approach to policy training, reminiscent of staged training strategies seen in prior works, by first freezing the low-level policy and only fine-tuning the language model. This preliminary phase focuses on aligning the embeddings produced by the MLLM with the feature space of the policy, to prevent the initial unstable gradient due to the mismatch between the pre-trained policy and the pre-trained LLM.
Stage 2: Fine-tuning end-to-end. In this stage, we unfreeze the low-level policy and proceed with fine-tuning the entire model in an end-to-end manner. This phase constitutes the majority of the training process.
@@
$\mathcal{L}=$ $\lambda_{1}\mathcal{L}_{\textnormal{policy}}(\pi_{\theta},o_{t},z_{{\texttt{<ACT>}}},a_{t})+\lambda_{2}\mathcal{L}_{\textnormal{LLM}}(f_{\phi},x_{\textnormal{txt}},x_{\textnormal{img}})$ (3)
$+\lambda_{3}\mathcal{L}_{\textnormal{CLIP}}(z_{{\texttt{<ACT>}}},g_{\textnormal{txt}}),$
where $\mathcal{L}_{\textnormal{policy}}(\cdot)$ is the diffusion denoise loss of low-level policy [16]. $\mathcal{L}_{\textnormal{LLM}}(\cdot)$ is the same auto-regressive training objective following [5], with a special token <ACT> instead. $\mathcal{L}_{\textnormal{CLIP}}(z_{\textnormal{act}},g_{\textnormal{txt}})$ is used to regularize the latent embedding $z_{{\texttt{<ACT>}}}$, ensuring it well aligned with the lower level ground truth text description $g_{txt}$ [39]. Specifically, we employ $\mathcal{L}_{\textnormal{CLIP}}(z_{\textnormal{act}},g_{\textnormal{txt}})=1-\cos(\textnormal{stop\_gradient}(\textnormal{clip}(g_{\textnormal{txt}})),z_{\textnormal{act}})$. This auxiliary loss helps regularize the predicted embedding. We set $\lambda_{1}=1$, $\lambda_{2}=100$, and $\lambda_{3}=1$. The $\lambda_{2}$ owns a high value because the LLM necessitates extensive finetuning due to the newly designed special tokens.
@@ V Experiments
Implementation Details. During training, we use LoRA with a rank of 16. All the reported results employ the pre-trained LLaVA-v1.5-7b [5] as VLM $f_{\phi}$ and CLIP-VIT-L/14-336 [53] as vision encoder $f_{enc}$. The implementation of low-level policy follows [16]. The fine-tuning data is a mixture of the CALVIN A, B, and C splits and the RAMA training set, combined in a ratio of 0.7 to 0.3, respectively. The fine-tuning of RationalVLA involves two stages: 5,000 iterations in the first stage and 30,000 in the second. The total training process takes 36 hours on eight 80GB H800 GPUs.
@@ V-A Baselines
Fine-tuned VLA model. With inherited language abilities, we fine-tune a typical VLA model, RoboFlamingo, to output no in response to defective instructions.

### v2

@@ Abstract
A fundamental requirement for real-world robotic deployment is the ability to understand and respond to natural language instructions. Existing language-conditioned manipulation tasks typically assume that instructions are perfectly aligned with the environment. This assumption limits robustness and generalization in realistic scenarios where instructions may be ambiguous, irrelevant, or infeasible. To address this problem, we introduce RAtional MAnipulation (RAMA), a new benchmark that challenges models with both unseen executable instructions and defective ones that should be rejected. In RAMA, we construct a dataset with over 14,000 samples, including diverse defective instructions spanning six dimensions: visual, physical, semantic, motion, safety, and out-of-context. We further propose the Rational Vision-Language-Action model (RationalVLA). It is a dual system for robotic arms that integrates the high-level vision-language model with the low-level manipulation policy by introducing learnable latent space embeddings. This design enables RationalVLA to reason over instructions, reject infeasible commands, and execute manipulation effectively. Experiments demonstrate that RationalVLA outperforms state-of-the-art baselines on RAMA by a 14.5% higher success rate and 0.94 average task length, while maintaining competitive performance on standard manipulation tasks. Real-world trials further validate its effectiveness and robustness in practical applications. Our project page is https://irpn-eai.github.io/RationalVLA/.
@@ I Introduction
Embodied intelligence represents the ultimate manifestation of artificial intelligence [1]. A necessary condition for the successful deployment of embodied intelligence in the real world is its ability to understand natural language and respond appropriately, either by providing correct answers or by executing the corresponding actions. This demand has sparked research on language-conditioned manipulation tasks, which require robots to follow natural language instructions to complete specific manipulation actions. Recent works have identified and investigated several key challenges in this field, particularly in handling complex and previously unseen instructions [2, 3, 4]. They primarily leverage the language generalization capabilities of pretrained vision-language models (VLMs) [5, 6, 7, 8, 9, 10], fine-tuning them on robotic datasets to obtain vision-language-action (VLA) models [11, 2, 12, 13, 14, 15], which enables the handling of various instructions to some extent.
@@ II Related Work
While previous research[21, 22, 23, 24] has made notable progress in enabling robots to perform precise and efficient manipulation, our work specifically focuses on the challenges of language-conditioned manipulation. To enhance the robustness of instruction understanding in such robots, prior research relevant to our proposed RationalVLA model can be organized into three key areas: (1) language-conditioned manipulation, which focuses on learning action policies from natural language. (2) dual-system VLA models, which decouple high-level reasoning from low-level control. and (3) efforts addressing defective instructions in human-robot collaboration of RAMA benchmark. Below, we briefly review representative work in each area.
@@ II-A Language-conditioned Manipulation
Early works [25, 26, 11] use text encoders to map language instructions to latent features, which are then conditioned by robot policies to predict actions. Some approaches [27, 28, 16, 29] employ diffusion models as action policies. RT-2 [2] and subsequent works [30, 31] integrate large language models to enhance language generalization. Inner Monologue [32] enriches instruction processing and planning in robotic control. RoboFlamingo [3] fine-tunes vision-language models for action prediction, while GR-1 [33] learns from action videos and Vision Language Planning [34] generate futural action videos from language to predict action. RoboMamba [35] utilizes state space models for efficient instruction-based reasoning. VLAS [36] extends the text language to speech. Our model demonstrates superior reasoning capabilities over these methods, enabled by its effective understanding of the intricate relationship between language and the environment. Another work [37] is a dual-arm robot framework combining dynamic manipulation primitives to efficiently unfold garments using self-supervised learning, achieving high coverage without expert data.
@@ II-C Defective Instructions in Human-robot Interaction
In human-robot interaction, accurate human instructions can significantly improve robotic task performance [46]. However, defective instructions can make robot confused. Ob-VLN [47] first considers the defective instructions in vision-language navigation. The discrepancies in actual scenes and given instructions can cause major failures. This work constructs a virtual graph to help the robot manage it. BadNaver [48] further identifies defective instructions as a type of jailbreak attack, which leads the model to execute vulnerable actions. Experiments on various robot agents imply that these instructions affect the safety. These works emphasize the importance of dealing with defective instructions. However, these works limit the question in navigation tasks, overlooking its importance in manipulation tasks.
@@ III-A Task Setting
  • Physical. The physical non-executable task refers to the target with non-existent shape and size in the scenario. Understanding these spatial positions sets requirements for the physical prior knowledge of the robot assistant.
@@
  • Motion. The task with non-executable motion refers to actions that can not be executed due to the limitations of the structure and kinematic characteristics of the mechanical arm. This requires the robot assistant to assess the feasibility of executing the action based on its capabilities.
@@
  • Out-of-context. In addition to the above dimensions, some completely unrelated, illogical, and nonsensical commands should also be included as non-executable instructions to ensure the robustness of the robot assistant against noise.
@@
Figure 2: In the RAMA benchmark, we assess models across six dimensions: visual, physical, semantic, motion, safety, and out-of-context.
@@ III-B Simulation Environment
To establish a unified and convenient benchmark, we construct a dataset and evaluation protocol based on CALVIN [17], which is well-regarded in language-conditioned manipulation research. 1) It ensures fair comparisons between different models and reproducibility of experiments. 2) It supports long-horizon tasks, which is crucial for evaluating the influence of defective instructions.
CALVIN is built on top of the PyBullet [51] simulator. It involves a desk with some objects and a Franka Panda Robot arm that manipulates the scene. It comprises 34 tasks across 4 distinct environments (A, B, C, and D), shown in Fig.3. The A, B, and C environments are used for training and D is used for testing to challengae the generalization. This setting is abbreviated as ABC→D in subsequent sections.
Figure 3: Experiment setting of CALVIN. CALVIN consists of four simulated environments (A, B, C, and D), which features different textures and object positions.
@@ III-C Dataset
We then convert the language annotations into the chat-like expression of MLLM assistants. CALVIN contains one language instruction and a list of (observation, action) pairs $[x_{\textnormal{task}},(o_{0},a_{0},o_{1},a_{1},...,o_{t},a_{t},...)]$ per trajectory. We programmatically generate texts in the format of chat interactions using templates. A simple example of user-assistant interactions is “User: Can you help me $x_{\textnormal{task}}$? Assistant: Sure, I will $x_{\textnormal{task}}$ <ACT>.” For defective instructions, it will be “User: Can you help me $x_{\textnormal{task}}$? Assistant: Sorry, I can not $x_{\textnormal{task}}$ <REJ>.” This process trains the model to recognize and appropriately respond to everyday human instructions, making informed decisions on whether to execute or refrain from executing actions. It thereby fosters a conversational interface capable of seamlessly transitioning from dialogue to actionable responses.
@@ III-D Evaluation
Figure 6: Overview of RationalVLA. Given the visual observation and text instruction that may be defective, an MLLM generates a text description of a task and an <ACT> token or a <REJ> token. The latent space embeddings from the <ACT> token’s last layer serve as a high-level latent goal for the downstream policy and those of <REJ> token reject defective instructions in the current scene. Our dual-system model can run the high-level MLLM reasoning and low-level action policy execution loops asynchronously. During inference, the action policy frequently updates actions based on environment changes and the latest <ACT> token’s embedding, while the MLLM updates less frequently, enabling efficient inference.
@@ IV-A Comparison of Architectures
Recent advances in VLA models have enabled various strategies for grounding instructions to robot actions. In this section, we compare three representative architectures, fine-tuned VLAs, VLM-based identifiers, and our proposed dual-system VLA.
Fine-tuned VLA. Given the strong generalization capabilities on visual and language modalities demonstrated by pretrained VLAs, a natural approach is to fine-tune an end-to-end VLA model to process and understand complex instructions. However, several key challenges limit its performance: (1) Fine-tuning a VLA model often leads to catastrophic forgetting [18], resulting in degraded language understanding capabilities, e.g., RoboFlamingo reaches a mediocre performance on the enriched instructions. (2) Considering downstream robotic tasks, VLA models frequently underperform specialized expert policies, e.g., RoboFlamingo lags behind 3DDA on standard CALVIN benchmarks.
VLM as Identifier. To address the aforementioned issues, a straightforward idea is to leverage both the multimodal reasoning capabilities of pretrained VLMs and the manipulation proficiency of a task-specific policy. Specifically, a separately trained VLM can serve as a high-level identifier that decides whether to reject a given instruction, after which a downstream policy is responsible for task execution. This constitutes a rudimentary dual-system architecture. With the capabilities of powerful VLMs such as GPT-4o, this design can effectively filter out inappropriate instructions. However, such a naive combination fails to fully exploit the VLM’s language understanding capabilities. When faced with complex instructions, the downstream policy often lacks sufficient comprehension, leading to task failure. Consequently, this approach falls short of achieving our overarching goal: aligning the robotic agent’s action space with the complexity of natural language instructions.
@@ IV-B Details of Our Dual-system VLA
We expand the vocabulary of the language model with an additional <ACT> token and a <REJ> token. The <ACT> token extracts the vision-language information, representing the latent space embedding of the action entity. Upon generating <ACT> token, its hidden embedding is further fed into the low-level policy as a condition. Some previous works have focused on dynamically processing neural network inputs [19, 20], which is equally effective for robots to flexibly handle language instructions. The model is trained to output either <ACT> tokens $y_{\textnormal{<ACT>}}$ or <REJ> tokens $y_{\textnormal{<REJ>}}$ depending on whether the instructions are executable or not in the current scene. Then, the $y_{\textnormal{<ACT>}}$ is projected by a projector $\varPsi$ and further unsqueezed in the zero dimension to align with the policy latent conditioning space. Thus, we get the latent embeddings $z=\mathsf{unsqueeze}(\varPsi(y_{\textnormal{<ACT>}}))$. For $y_{\textnormal{<REJ>}}$, it is projected to be $z_{0}=\varPsi(y_{\textnormal{<REJ>}})$, which is an all-zero tensor to stop action. Finally, $z$ is fed into the policy $\pi_{\theta}$.
@@ IV-C Training
The training of RationalVLA employs a dual-system technique to integrate the MLLM and policy components. We leverage Low Rank Adaptation [52] (LoRA) for the supervised fine-tuning of the MLLM, allowing for more efficient training.
Stage 1: Pre-training for feature alignment. We adopt a cold start approach to policy training, reminiscent of staged training strategies seen in prior works, by first freezing the low-level policy and only fine-tuning the language model. This preliminary phase focuses on aligning the embeddings produced by the MLLM with the feature space of the policy to prevent the initial unstable gradient due to the mismatch between the pre-trained policy and the pre-trained LLM.
Stage 2: Fine-tuning in an end-to-end manner. In this stage, we unfreeze the low-level policy and proceed with fine-tuning the entire model in an end-to-end manner. This phase constitutes the majority of the training process.
@@
$\mathcal{L}=$ $\lambda_{1}\mathcal{L}_{\textnormal{policy}}(\pi_{\theta},o_{t},z_{\textnormal{{{<ACT>}}}},a)+\lambda_{2}\mathcal{L}_{\textnormal{LLM}}(f_{\phi},x_{\textnormal{txt}},x_{\textnormal{img}})$ (3)
$+\lambda_{3}\mathcal{L}_{\textnormal{CLIP}}(z_{\textnormal{{{<ACT>}}}},g_{\textnormal{txt}}),$
where $\mathcal{L}_{\textnormal{policy}}(\cdot)$ is the diffusion denoise loss of low-level policy [16]. $\mathcal{L}_{\textnormal{LLM}}(\cdot)$ is the same auto-regressive training objective following [5], with a special token <ACT> instead. $\mathcal{L}_{\textnormal{CLIP}}(z_{\textnormal{act}},g_{\textnormal{txt}})$ is used to regularize the latent embedding $z_{\textnormal{{{<ACT>}}}}$, ensuring it well aligned with the lower level ground truth text description $g_{\textnormal{txt}}$ [38]. Specifically, we employ $\mathcal{L}_{\textnormal{CLIP}}(z_{\textnormal{act}},g_{\textnormal{txt}})=1-\cos(\textnormal{stop\_gradient}(\textnormal{clip}(g_{\textnormal{txt}})),z_{\textnormal{act}})$. This auxiliary loss helps regularize the predicted embedding. We set $\lambda_{1}=1$, $\lambda_{2}=100$, and $\lambda_{3}=1$. The $\lambda_{2}$ owns a high value because the LLM necessitates extensive finetuning due to the newly designed special tokens.
@@ V Experiments
Implementation Details. During training, we use LoRA with a rank of 16. All the reported results employ the pre-trained LLaVA-v1.5-7b [5] as VLM $f_{\phi}$ and CLIP-VIT-L/14-336 [53] as vision encoder $f_{\textnormal{encoder}}$. The implementation of low-level policy follows [16]. The fine-tuning data is a mixture of the CALVIN A, B, and C splits and the RAMA training set, combined in a ratio of 0.7 to 0.3, respectively. The fine-tuning of RationalVLA involves two stages: 5,000 iterations in the first stage and 30,000 in the second. The total training process takes 36 hours on eight 80GB H800 GPUs.
@@ V-A Baselines
Fine-tuned VLA model. With inherited language abilities, we fine-tune a typical VLA model, RoboFlamingo [3], to do not output action in response to defective instructions.